# ECHO: Efficient Dataset Condensation by Higher-Order Distribution Alignment

### 26 Dec 2023 v1

@@ Abstract
In the era of deep learning, training deep neural networks often requires extensive data, leading to substantial costs. Dataset condensation addresses this by learning a small synthetic set that preserves essential information from the original large-scale dataset. Nowadays, optimization-oriented methods dominate dataset condensation for state-of-the-art (SOTA) results, but their computationally intensive bi-level optimization hinders practicality with large datasets. To enhance efficiency, as alternative solutions, Distribution-Matching (DM)-based methods reduce costs by aligning the representation distributions of real and synthetic examples. However, current DM-based methods still yield less comparable results to SOTA optimization-oriented methods. In this paper, we argue that existing DM-based methods overlook the higher-order alignment of the distributions, which may lead to sub-optimal matching results. Inspired by this, we propose a new DM-based method named as Efficient Dataset Condensation by Higher-Order Distribution Alignment (ECHO). Specifically, rather than only aligning the first-order moment of the representation distributions as previous methods, we learn synthetic examples via further aligning the higher-order moments of the representation distributions of real and synthetic examples based on the classical theory of reproducing kernel Hilbert space. Experiments demonstrate the proposed method achieves a significant performance boost while maintaining efficiency across various scenarios.
1 Introduction
Recently, the utilization of large-scale datasets comprising millions or even billions of examples has become an indispensable prerequisite for achieving state-of-the-art (SOTA) models (Zhao and Bilen 2021a; Xia et al. 2022; Li et al. 2022a, 2023a, b, 2023b; Zhang et al. 2024). However, the associated storage expenses and computational costs involved in training these models present formidable challenges, often rendering them beyond the reach of startups and non-profit organizations (Wang et al. 2018; Coleman et al. 2019; Sorscher et al. 2022; Pooladzandi, Davini, and Mirzasoleiman 2022; Zheng et al. 2022; Jin et al. 2022; Yang et al. 2022).
@@ 1 Introduction
To alleviate the challenges associated with larger datasets, Dataset Condensation (DC) (Wang et al. 2018; Yu, Liu, and Wang 2023; Lei and Tao 2023; Liu et al. 2022; Loo et al. 2022; Cui et al. 2022; Wang et al. 2023) has emerged to reduce the training cost by synthesizing a compact set of informative examples. Since its proposal, DC has attracted significant attention for addressing the challenges posed by the data burden (Cazenavette et al. 2022; Zhao and Bilen 2021b; Kim et al. 2022; Wang et al. 2022; Zhao and Bilen 2023; Liu et al. 2023a). Typically, DC methods condense the dataset by minimizing the distance between real and synthetic examples via a pre-defined matching objective. As various matching objectives can lead to different optimization processes, based on whether to perform a costly bi-level optimization (Liu et al. 2021), these methods can mainly be categorized into two groups (Zhao et al. 2023a): (1) Optimization-oriented methods (Wang et al. 2018; Zhao and Bilen 2021b; Kim et al. 2022; Zhao and Bilen 2021a; Cazenavette et al. 2022), which usually generate condensed examples by conducting performance matching or parameter matching (Yu, Liu, and Wang 2023) via a bi-level optimization; (2) Distribution-Matching (DM)-based methods (Wang et al. 2022; Zhao and Bilen 2023; Zhao et al. 2023a; Sajedi et al. 2023), which focus on aligning the representation distributions between real and synthetic data. Optimization-oriented methods have faced criticism for their inefficiency, primarily due to the involvement of bi-level optimization modules and time-consuming network updating processes (Zhang et al. 2023; Wang et al. 2022; Zhao and Bilen 2023). In contrast, DM-based methods do not involve such nested optimization of models, which significantly reduces the computational cost associated with dataset condensation. Nevertheless, the informativeness of the condensed examples generated by current DM-based methods may not be as comparable to those produced by optimization-oriented methods.
In this paper, we argue that the higher-order alignment of distributions is important for achieving good distribution matching, while existing DM-based methods neglect this and may result in sub-optimal results. As illustrated in Fig. 1, despite sharing the same first moment, the representation distributions of original and synthetic examples with misaligned second-order moments (Fig. 0(a)) or third-order moments (Fig. 0(b)) can exhibit much distinct characteristics. Since previous DM-based methods (Wang et al. 2022; Zhao and Bilen 2023; Zhao et al. 2023a; Sajedi et al. 2023) only align the first-order moment of the representation distributions, their objectives may lead to much different distributions with the misalignment of higher-order moments.
Motivated by this issue, we propose a new DM-based method named as Efficient Dataset Condensation by Higher-Order Distribution Alignment (ECHO). Specifically, unlike previous DM-based methods that only aligning the first-order moment of the representation distributions, we learn synthetic examples via further aligning the higher-order moments of the representation distributions of real and synthetic examples by transforming the distributions into a Reproducing Kernel Hilbert Space (RKHS). Based on the classical kernel theory (Muandet et al. 2017; Altun and Smola 2006; Borgwardt et al. 2006), this transformation allows us to represent an infinite number of moments in a kernel-function form. After that, by leveraging empirical estimation, we can readily align the infinite order of moments of the representation distributions via minimizing the distance between the kernel embeddings of the real and synthetic data, which is also known as Maximum Mean Discrepancy (MMD) (Altun and Smola 2006). Note that the DM method (Zhao and Bilen 2023) claimed its method utilizes the empirical MMD to match the distributions, but the randomly initialized models are not valid embedding functions for MMD (Zhao et al. 2023b) and thus it has no theoretical guarantee, while our ECHO exploits the classical kernel method to address the misalignment issue of higher-order moments of the representation distributions with theoretical guarantees.
Comprehensive experiments across various datasets show that our ECHO method can efficiently generate highly informative condensed examples, which not only maintains the efficiency as the original DM method, but also exhibits significant improvements. Remarkably, the efficiency of our method makes it easily applicable to realistic and larger datasets like ImageNet (Deng et al. 2009). As shown in Section 4.2, our ECHO can achieve superior results compared to the current SOTA optimization-oriented method IDC (Kim et al. 2022) in the high-resolution ImageNet subsets, while requiring considerably less training time (around 30× $\sim$ 64x speed up).
@@
  • We reveal and analyze the importance of the alignment of higher-order moments for distribution matching, which is overlooked by previous methods.
  • We propose a theoretical-guaranteed method for dataset condensation named ECHO, which applies the classical kernel method to represent an infinite number of moments in a kernel-function form, enabling the improved alignment of the higher-order moments of the representation distributions.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of the proposed method, where ECHO yields excellent performance across various scenarios, showing the applicability to high-resolution data, the generalization for cross-architecture training, and the practicality to continual learning task.
2 Background
@@
where $\bm{D}$ represents a distance metric such as Mean Square Error (MSE), and $\phi$ denotes the matching objective. As mentioned before, various objectives can lead to different optimization processes (Yu, Liu, and Wang 2023), and based on whether to perform a costly bi-level optimization, existing methods can be mainly divided into optimization-oriented methods and Distribution-Matching (DM)-based methods.  [^1] Note that the introduction about more dataset condensation works can be found in Appendix B.
@@
which works as minimizing the gap between empirical first moment of the representation distributions between $\mathcal{S}$ and $\mathcal{T}$. Compared to previous optimization-oriented methods, DM (Zhao and Bilen 2023) eliminates the need for network updating, relying instead on randomly initialized encoders. As a result, the costly bi-level optimization is avoided, leading to significantly improved training efficiency.
@@
3 Methodology
In this section, we begin by analyzing the importance of the alignment of higher-order moments for distribution matching. Subsequently, we propose our method ECHO by exploiting the classical kernel method (Altun and Smola 2006; Borgwardt et al. 2006) to align the higher-order moments of the representation distributions between real and synthesized data with theoretical guarantees.
3.1 Importance of the Higher-Order Alignment
First of all, as shown in Eq. (3), it is evident that DM (Zhao and Bilen 2023) only considers aligning the first moment (mean) of the representation distributions, while neglecting higher-order moments. At a high level, it may lead to the higher-order misalignment of the representation distributions of its condensed data and original data.
Then, to verify this misalignment issue and showcase the importance of the higher-order alignment, we evaluate the distance of the moments between the condensed set and the original training set by adding the regularization of aligning the higher-order moment to DM on CIFAR10 with the number of images per class being 10. The results are shown in Tab. 1. As can be seen, by adding the second-order regularization, the distance between higher-order moments of the obtained data and original data can be significantly reduced, which demonstrates only aligning the first moment will lead to the higher-order misalignment of the representation distributions. In addition, by performing more regularization, the performance of the condensed dataset will be further improved with the increasing higher-order alignment. It clearly shows that the alignment of higher-order moments of representation distributions is important for distribution matching, which is overlooked by previous works.
Table 1: The distance between the moments of the condensed set and the original training set. “+2nd(3rd)-order Reg.” denotes adding the regularization of aligning the 2nd(3rd)-order moment to the original loss of DM (Zhao and Bilen 2023).
@@ 3.1 Importance of the Higher-Order Alignment
+2nd-order Reg. 4.51 6.96 6.35 52.1 ($\uparrow$ 3.2)
+2nd- and 3rd-order Reg. 3.69 6.14 5.94 53.9 ($\uparrow$ 5.0)
ECHO 0.82 1.23 1.64 63.5 ($\uparrow$ 14.6)
3.2 Efficient Dataset Condensation by Higher-Order Distribution Alignment
From the preceding analysis, it becomes evident that perfecting distribution matching necessitates the consideration of higher-order moments. While incorporating higher-order regularizations directly aids in aligning these moments, it is limited to finite moments. Moreover, tuning the regularization coefficient becomes increasingly challenging with a growing number of regularization terms. In this subsection, we represent a new DM-based method that aligns the infinite order of moments in a kernel-function form. We depict the framework of the proposed ECHO in Fig. 2.
Figure 2: The framework of ECHO, after extracting the representations via a encoder network, the distributions of real and synthetic representations are further embedded in the Reproducing Kernel Hilbert Space (RKHS), where the ECHO loss $\mathcal{L}_{\text{ECHO}}=\hat{\text{MMD}}^{2}(\bm{P}_{\mathcal{T}},\bm{P}_{\mathcal{S}})$ is calculated to guild the update of synthetic examples for higher-order distribution alignment.
@@ 3.2 Efficient Dataset Condensation by Higher-Order Distribution Alignment
where $\mathcal{K}_{X,Y}=E_{X,Y}[\mathcal{K}(g_{\theta}(x),g_{\theta}(y))]$ with $x\sim X,y\sim Y$. Due to limited page, we provide the derivation of Eq. (8) in Appendix C. Last, note that we only have access to the datasets $\mathcal{T}$ and $\mathcal{S}$ rather than their underlying distributions. In order to tackle this issue, denoting the empirical approximation of $\mu[\bm{P}_{\mathcal{T}}]$ and $\mu[\bm{P}_{\mathcal{S}}]$ as $\mu[\mathcal{T}]=\frac{1}{|\mathcal{T}|}\sum_{i=1}^{|\mathcal{T}|}\mathcal{K}(g_{\theta}(\bm{x}_{i}),\cdot)$, $\mu[\mathcal{S}]=\frac{1}{|\mathcal{S}|}\sum_{j=1}^{|\mathcal{S}|}\mathcal{K}(g_{\theta}(\bm{s}_{j}),\cdot)$ respectively, we introduce the following theorem:
@@
Theorem 2 guarantees that the empirical approximations $\mu[\mathcal{T/S}]$ are good proxies for $\mu[\bm{P}_{\mathcal{T/S}}]$. Therefore, we can modify Eq. (8) to the following empirical form as the ECHO loss:
$\mathcal{L}_{\text{ECHO}}=\hat{\text{MMD}}^{2}(\bm{P}_{\mathcal{T}},\bm{P}_{\mathcal{S}})=\hat{\mathcal{K}}_{\mathcal{T,T}}+\hat{\mathcal{K}}_{\mathcal{S,S}}-2\hat{\mathcal{K}}_{\mathcal{T,S}},$ (9)
where $\hat{\mathcal{K}}_{X,Y}=\frac{1}{|X|\cdot|Y|}\sum_{i=1}^{|X|}\sum_{j=1}^{|Y|}\mathcal{K}(g_{\theta}(x_{i}),g_{\theta}(y_{j}))$ with $\{x_{i}\}_{i=1}^{|X|}\sim X,\{y_{j}\}_{j=1}^{|Y|}\sim Y.$ Based on the analysis above, we have successfully achieved the transformation of an infinite number of moments into a finite form using RKHS. As shown in Tab. 1, this transformation allows us to effectively align the distributions between $\mathcal{T}$ and $\mathcal{S}$ during the condensing process.
3.3 Training Algorithm of ECHO
The pseudo-code of ECHO is provided in Appendix A. In addition to the kernel method, we exploit the following two techniques to enhance the distribution matching.
@@ 3.3 Training Algorithm of ECHO
Random Herding K-Center DC DSA CAFE CAFE+DSA DM KIP MTT IDM ECHO
@@
In this way, the storage space of $\mathcal{S}$ can be further leveraged. Following IDC (Kim et al. 2022) and IDM (Zhao et al. 2023a), the same factor technique is incorporated into our framework, where we further exploit its benefits in aligning distributions in higher-order moments.
Iteration per Random Model. Following DM (Zhao and Bilen 2023), we employ multiple randomly initialized models to extract representation embeddings from both $\mathcal{T}$ and $\mathcal{S}$. In contrast to DM, where only a single-step iteration is performed for each model, we posit that relying solely on the representation distributions of one batch of real and synthetic examples may introduce biased matching targets. To address this, without incurring additional memory usage, we empirically observe that conducting multiple iterations per model (IPM) enhances the performance of the condensed set.
4 Experiments
Table 3: The performance and efficiency comparison on high-resolusion ImageNet-subsets. The synthetic examples are condensed using ResNetAP-10. The minimal time required for obtaining the best performance is reported, which is measured on a single RTX-A6000 GPU with same batch size. Following previous works, for ImageNet-100, all methods are splitted into five sub-tasks with 20 classes each for faster optimization.
@@ 4 Experiments
ECHO 73.4 1.1h 76.8 3.1h 46.9 3.5h 55.5 4.2h
In this section, we assess the efficacy of ECHO in enhancing both the performance of dataset condensation and the efficiency. We introduce the experimental setups for reproducibility (Section 4.1), compare our method with SOTA benchmarks, present the efficiency evaluations and generalization results across unseen architectures (Section 4.2), and conduct comprehensive ablation studies (Section 4.3). More analysis and visualizations are presented in the Appendix.
4.1 Experimental Setups
Datasets. We evaluate the classification performance of networks trained on synthetic examples that have been condensed using various baselines as well as our proposed ECHO method. Our evaluation encompasses four low-resolution datasets: F-MNIST (Xiao, Rasul, and Vollgraf 2017), SVHN (Netzer et al. 2011), CIFAR-10 (Krizhevsky, Hinton et al. 2009), and CIFAR-100 (Krizhevsky, Hinton et al. 2009). In addition, we also conduct experiments on the high-resolution ImageNet subsets (Deng et al. 2009). Detailed descriptions of datasets can be found in Appendix D.
@@ 4.1 Experimental Setups
Baselines. We employ an extensive range of methods as baselines for comparison. Regarding coreset selection methods, we consider the following: (1) Random, (2) Herding (Welling 2009), and (3) K-Center (Farahani and Hekmatfar 2009; Sener and Savarese 2017). For optimization-oriented DC methods, we evaluate (4) DC (Zhao and Bilen 2021b), (5) DSA (Zhao and Bilen 2021a), (6) IDC (Kim et al. 2022), (7) KIP (Nguyen et al. 2021) and (8) MTT (Cazenavette et al. 2022). On the other hand, for DM-based DC methods, we include (9) CAFE (Wang et al. 2022), (10) its variant CAFE+DSA (Wang et al. 2022), (11) DM (Zhao and Bilen 2023) and (12) IDM (Zhao et al. 2023a). We provide detailed descriptions of baselines in Appendix E.
@@
4.2 Comparison to the SOTA Methods
@@
To further demonstrate the advantages of our method, we provide the test accuracy across varying training steps in Fig. 3. As observed, our method consistently outperforms DM (Zhao and Bilen 2023) at different training steps. Even without the factor technique, our method still achieves considerable improvement, highlighting the effectiveness of ECHO in aligning distributions compared to DM.
Figure 3: Performance comparison between ECHO and DM across varying training steps. ECHO w/o Fac. denotes the ECHO without using the factor technique.
Cross-Architecture Evaluation.  We further assess the performance of our condensed examples on different architectures. In Tab. 4, we present the performance of our condensed examples from the CIFAR-10 dataset on ConNet-3, ResNet-10 (He et al. 2016), and DenseNet-121 (Huang et al. 2017). Combining the results from Tab. 2, we can find that ECHO outperforms the compared methods not only on the architecture used for condensation but also on unseen architectures.

### 25 Feb 2024 v3

# M3D: Dataset Condensation by Minimizing Maximum Mean Discrepancy

@@ Abstract
Training state-of-the-art (SOTA) deep models often requires extensive data, resulting in substantial training and storage costs. To address these challenges, dataset condensation has been developed to learn a small synthetic set that preserves essential information from the original large-scale dataset. Nowadays, optimization-oriented methods have been the primary method in the field of dataset condensation for achieving SOTA results. However, the bi-level optimization process hinders the practical application of such methods to realistic and larger datasets. To enhance condensation efficiency, previous works proposed Distribution-Matching (DM) as an alternative, which significantly reduces the condensation cost. Nonetheless, current DM-based methods still yield less comparable results to SOTA optimization-oriented methods. In this paper, we argue that existing DM-based methods overlook the higher-order alignment of the distributions, which may lead to sub-optimal matching results. Inspired by this, we present a novel DM-based method named M3D for dataset condensation by Minimizing the Maximum Mean Discrepancy between feature representations of the synthetic and real images. By embedding their distributions in a reproducing kernel Hilbert space, we align all orders of moments of the distributions of real and synthetic images, resulting in a more generalized condensed set. Notably, our method even surpasses the SOTA optimization-oriented method IDC on the high-resolution ImageNet dataset. Extensive analysis is conducted to verify the effectiveness of the proposed method. Source codes are available at https://github.com/Hansong-Zhang/M3D.
Introduction
In the era of deep learning, the utilization of large-scale datasets comprising millions of samples has become an indispensable prerequisite for achieving state-of-the-art (SOTA) models (Zhao and Bilen 2021a; Xia et al. 2022). However, the associated storage expenses and computational costs involved in training these models present formidable challenges, often rendering them beyond the reach of startups and non-profit organizations (Wang et al. 2018; Coleman et al. 2019; Sorscher et al. 2022).
@@ Introduction
To alleviate the challenges associated with larger datasets, Dataset Condensation (DC) (Wang et al. 2018) has emerged to reduce the training cost by synthesizing a compact set of informative images. Since its proposal, DC has attracted significant attention for addressing the challenges posed by the data burden (Cazenavette et al. 2022; Zhao and Bilen 2021b; Kim et al. 2022; Wang et al. 2022; Zhao and Bilen 2023). Typically, DC condenses the dataset by minimizing the distance between real and synthetic images via a pre-defined metric. Based on whether to perform a costly bi-level optimization (Liu et al. 2021), these methods can be generally categorized into two groups: (1) Optimization-Oriented methods (Zhao and Bilen 2021b; Kim et al. 2022; Zhao and Bilen 2021a; Cazenavette et al. 2022), which usually generate condensed examples by conducting performance matching or parameter matching via a bi-level optimization (Yu, Liu, and Wang 2023); (2) Distribution-Matching(DM)-based methods (Wang et al. 2022; Zhao and Bilen 2023), which focus on aligning the feature distributions between real and synthetic data. Optimization-oriented methods have faced criticism for their inefficiency, primarily due to the involvement of bi-level optimization modules and time-consuming network updating processes (Zhang et al. 2023; Wang et al. 2022; Zhao and Bilen 2023). In contrast, DM-based methods do not involve such nested optimization of models, which significantly reduces the computational cost associated with dataset condensation. Nevertheless, the informativeness of the condensed examples generated by current DM-based methods may not be as comparable to those produced by optimization-oriented methods.
In this paper, we address a crucial oversight in existing DM-based methods, which is their neglect of higher-order moments of the distribution. As illustrated in Fig. 1, despite sharing the same first moment, the representation distributions of original and synthetic examples with misaligned second-order moments (Fig. 0(a)) or third-order moments (Fig. 0(b)) can exhibit much distinct characteristics. Motivated by this issue, we propose a novel DM-based method involving Minimizing the Maximum Mean Discrepancy (M3D) between the representation distributions of the real and synthetic images. Unlike previous DM-based methods that solely embed images in a feature representation space and align the first moment, our method further embeds the distribution of feature representations into a reproducing kernel Hilbert space. This transformation allows us to represent the infinite order of moments in a kernel-function form. By leveraging empirical estimation, we can readily align both first- and higher-order moments of the real and synthetic data with theoretical guarantees. Our method not only maintains the efficiency of the DM-based method but also exhibits significant improvements. Remarkably, the efficiency of our method makes it easily applicable to realistic and larger datasets like ImageNet (Deng et al. 2009).
@@
  • We reveal the importance of the alignment of higher-order moments for distribution matching, which is overlooked by previous DM-based methods.
  • We propose a theoretical-guaranteed method for dataset condensation named M3D, which applies the classical kernel method to represent an infinite number of moments in a kernel-function form, enabling the improved alignment of the higher-order moments of the representation distributions.
  • We conduct extensive experiments to demonstrate the effectiveness and efficiency of our proposed method, where M3D yields SOTA performance with strong generalization across various scenarios.
Background
@@
where $\bm{D}$ represents a distance metric such as Mean Square Error (MSE), and $\phi$ denotes the matching objective. As mentioned before, various objectives can lead to different optimization processes (Yu, Liu, and Wang 2023), and based on whether to perform a costly bi-level optimization, existing methods can be mainly divided into optimization-oriented methods and Distribution-Matching (DM)-based methods.  [^1] Note that the introduction about more dataset condensation works can be found in the Appendix.
@@
which works as minimizing the gap between empirical first moment of the representation distributions between $\mathcal{S}$ and $\mathcal{T}$. Compared to previous optimization-oriented methods, DM (Zhao and Bilen 2023) eliminates the need for network updating, relying instead on randomly initialized encoders. Furthermore, the costly bi-level optimization is avoided in DM, leading to significantly improved training efficiency.
@@
Methodology
In this section, we begin by analyzing the importance of the alignment of higher-order moments for distribution matching. Subsequently, we propose our method M3D by exploiting the classical kernel method (Altun and Smola 2006; Borgwardt et al. 2006) to align the higher-order moments of the representation distributions between real and synthesized data with theoretical guarantees.
Importance of the Higher-Order Alignment
As shown in Eq. (3), it is evident that DM (Zhao and Bilen 2023) only considers aligning the first moment (mean) of the representation distributions, while neglecting higher-order moments. At a high level, it may lead to the higher-order misalignment of the representation distributions of its condensed data and original data.
To investigate this misalignment issue and highlight the importance of the higher-order alignment, we assessed the moment distances between the condensed set and the original training set on CIFAR-10 with 10 images per class. This was done by incorporating higher-order moment regularization terms into the original loss of DM (Zhao and Bilen 2023). The results, presented in Table 1, reveal that adding second-order regularization notably decreases the distance between higher-order moments of the condensed and original data, underscoring the inadequacy of aligning only the first moment. Furthermore, performing more regularization enhances the condensed dataset’s performance through improved higher-order alignment. These results underscore the critical role of higher-order moment alignment in distribution matching, which is neglected in previous works.
@@
+2nd Reg. 4.51 6.96 6.35 52.1 ($\uparrow$ 3.2)
+2nd &\&& 3rd Reg. 3.69 6.14 5.94 53.9 ($\uparrow$ 5.0)
M3D 0.82 1.23 1.64 63.5 ($\uparrow$ 14.6)
Table 1: The distance between the moments of the condensed set and the original training set. “+2nd(3rd) Reg.” denotes adding the regularization of aligning the 2nd(3rd)-order moment to the original loss of DM.
Minimizing Maximum Mean Discrepancy
From the preceding analysis, it becomes evident that perfecting distribution matching necessitates the consideration of higher-order moments. While incorporating higher-order regularizations directly aids in aligning these moments, it is limited to finite moments. Moreover, tuning the regularization coefficient becomes increasingly challenging with a growing number of regularization terms. In this subsection, we represent a new DM-based method that aligns the infinite order of moments in a kernel-function form. We depict the framework of the proposed M3D in Fig. 2.
Figure 2: The framework of M3D. After extracting the representations via a encoder network, the distributions of real and synthetic representations are further embedded in the Reproducing Kernel Hilbert Space (RKHS), where the M3D loss $\mathcal{L}_{\text{M3D}}$ is calculated to guild the update of synthetic examples for higher-order distribution alignment.
@@
where $\mathcal{K}_{X,Y}=E_{X,Y}[\mathcal{K}(g_{\theta}(x),g_{\theta}(y))]$ with $x\sim X,y\sim Y$. Due to limited page, we provide the derivation of Eq. (8) in the Appendix. Last, note that we only have access to the datasets $\mathcal{T}$ and $\mathcal{S}$ rather than their underlying distributions. In order to tackle this issue, denoting the empirical approximation of $\mu[\bm{P}_{\mathcal{T}}]$ and $\mu[\bm{P}_{\mathcal{S}}]$ as $\mu[\mathcal{T}]=\frac{1}{|\mathcal{T}|}\sum_{i=1}^{|\mathcal{T}|}\mathcal{K}(g_{\theta}(\bm{x}_{i}),\cdot)$, $\mu[\mathcal{S}]=\frac{1}{|\mathcal{S}|}\sum_{j=1}^{|\mathcal{S}|}\mathcal{K}(g_{\theta}(\bm{s}_{j}),\cdot)$ respectively, we introduce the following theorem:
@@
Theorem 2 guarantees that the empirical approximations $\mu[\mathcal{T/S}]$ are good proxies for $\mu[\bm{P}_{\mathcal{T/S}}]$. Therefore, we can modify Eq. (8) to the following empirical form as the M3D loss:
$\mathcal{L}_{\text{M3D}}=\hat{\text{MMD}}^{2}(\bm{P}_{\mathcal{T}},\bm{P}_{\mathcal{S}})=\hat{\mathcal{K}}_{\mathcal{T,T}}+\hat{\mathcal{K}}_{\mathcal{S,S}}-2\hat{\mathcal{K}}_{\mathcal{T,S}},$ (9)
where $\hat{\mathcal{K}}_{X,Y}=\frac{1}{|X|\cdot|Y|}\sum_{i=1}^{|X|}\sum_{j=1}^{|Y|}\mathcal{K}(g_{\theta}(x_{i}),g_{\theta}(y_{j}))$ with $\{x_{i}\}_{i=1}^{|X|}\sim X,\{y_{j}\}_{j=1}^{|Y|}\sim Y.$ Based on the analysis above, we have successfully achieved the transformation of an infinite number of moments into a finite form using RKHS. As shown in Table 1, this transformation allows us to effectively align the distributions between $\mathcal{T}$ and $\mathcal{S}$ during the condensing process.
Training Algorithm of M3D
The pseudo-code of M3D is provided in the Appendix. In addition to the kernel method, we exploit the following two techniques to enhance the distribution matching.
@@
In this way, the storage space of $\mathcal{S}$ can be further leveraged. Following previous works, the same factor technique is incorporated into our framework, where we further exploit its benefits in aligning distributions in higher-order moments.
Iteration per Random Model. Following DM (Zhao and Bilen 2023), we employ multiple randomly initialized models to extract representation embeddings from both $\mathcal{T}$ and $\mathcal{S}$. In contrast to DM, where only a single-step iteration is performed for each model, we posit that relying solely on the representation distributions of one batch of real and synthetic examples may introduce matching biases. To address this, without incurring additional memory usage, we empirically observe that conducting multiple iterations per model (IPM) enhances the performance of the condensed set.
Experiments
Figure 3: Performance comparison between M3D and DM across varying training steps. M3D w/o Fac denotes the M3D without using the factor technique.
@@
Random Herding K-Center DC DSA CAFE CAFE+DSA DM IDM M3D
MNIST 1 0.017 64.9±${}_{\pm 3.5}$ 89.2±${}_{\pm 1.6}$ 89.3±${}_{\pm 1.5}$ 91.7±${}_{\pm 0.5}$ 88.7±${}_{\pm 0.6}$ 93.1±${}_{\pm 0.3}$ 90.8±${}_{\pm 0.5}$ 89.7±${}_{\pm 0.6}$ 94.4±${}_{\pm 0.2}$ 99.6±${}_{\pm 0.0}$
10 0.17 95.1±${}_{\pm 0.9}$ 93.7±${}_{\pm 0.3}$ 84.4±${}_{\pm 1.7}$ 97.4±${}_{\pm 0.2}$ 97.8±${}_{\pm 0.1}$ 97.2±${}_{\pm 0.2}$ 97.5±${}_{\pm 0.1}$ 97.5±${}_{\pm 0.1}$ 97.6±${}_{\pm 0.1}$
50 0.83 97.9±${}_{\pm 0.2}$ 94.8±${}_{\pm 0.2}$ 97.4±${}_{\pm 0.3}$ 98.8±${}_{\pm 0.2}$ 99.2±${}_{\pm 0.1}$ 98.6±${}_{\pm 0.2}$ 98.9±${}_{\pm 0.2}$ 98.6±${}_{\pm 0.1}$ 98.2±${}_{\pm 0.2}$
@@
M3D 73.4 1.1h 76.8 3.1h 46.9 3.5h 55.5 4.2h
Table 3: The performance and efficiency comparison on high-resolusion ImageNet-subsets. The synthetic examples are condensed using ResNetAP-10. The minimal time required for obtaining the best performance is reported, which is measured on a single RTX-A6000 GPU with same batch size. For ImageNet-100, all methods are splitted into five sub-tasks with 20 classes each for faster optimization.
(a) Initialized SVHN images.
(b) Condensed images by DM.
(c) Condensed images by M3D.
Figure 4: Visualization of the condensed set of SVHN dataset with 10 images per class. The condensed set is generated by (b) DM and (c) M3D. Both DM and M3D use the same initialization as (a) shows.
In this section, we begin by comparing our proposed M3D with SOTA baselines on multiple benchmark datasets. Subsequently, we conduct an in-depth examination of M3D through ablation analysis.
Experimental Setups
Datasets. We evaluate the classification performance of networks trained on synthetic images that have been condensed using various baselines as well as our proposed method M3D. Our evaluation encompasses five low-resolution datasets: MNIST (LeCun et al. 1998), Fashion-MNIST (F-MNIST) (Xiao, Rasul, and Vollgraf 2017), SVHN (Netzer et al. 2011), CIFAR-10 (Krizhevsky, Hinton et al. 2009), and CIFAR-100 (Krizhevsky, Hinton et al. 2009). In addition, we also conduct experiments on the high-resolution dataset ImageNet subsets (Deng et al. 2009). Detailed descriptions of datasets can be found in the Appendix.
@@
Baselines. We employ an extensive range of methods as baselines for comparison. Regarding coreset selection methods, we consider the following: (1) Random, (2) Herding (Welling 2009), and (3) K-Center (Farahani and Hekmatfar 2009; Sener and Savarese 2017). For optimization-oriented DC methods, we evaluate (4) DC (Zhao and Bilen 2021b), (5) DSA (Zhao and Bilen 2021a), (6) IDC (Kim et al. 2022). On the other hand, for DM-based DC methods, we include (7) CAFE (Wang et al. 2022), (8) its variant CAFE+DSA (Wang et al. 2022), (9) DM (Zhao and Bilen 2023) and (10) IDM (Zhao et al. 2023a). We provide detailed descriptions of baselines in the Appendix.
@@
Comparison to the SOTA Methods
@@
To further demonstrate the advantages of our method, we provide the test accuracy across varying training steps in Fig. 3. As observed, our method consistently outperforms DM at different training steps. Even without the factor technique, our method still achieves considerable improvement, highlighting the effectiveness of M3D in aligning distributions compared to previous DM-based methods.
Cross-Architecture Evaluation.  We further assess the performance of our condensed examples on different architectures. In Table 4, we present the performance of our condensed examples from CIFAR-10 dataset on ConNet-3, ResNet-10 (He et al. 2016), and DenseNet-121 (Huang et al. 2017). Combining the results from Table 2, we can find that M3D outperforms the compared methods not only on the architecture used for condensation but on unseen ones.
Visualizations. We visualize the condensed images of SVHN and ImageNet in Fig. 4 and Fig. 5, respectively. For SVHN, we initialize the synthetic set $\mathcal{S}$ using random images from the training set $\mathcal{T}$ and then apply the condensation process using DM and M3D. As shown, the condensed images by DM and M3D appear as if the original images have been augmented with a distinct texture. Notably, the condensed images produced by our method exhibit a more pronounced and visually appealing texture compared to DM. While the overall appearance remains similar, our condensed images demonstrate better alignment with the higher-order moments of the original training set. In the case of ImageNet, the condensed images exhibit a texture reminiscent of a sunspot. In contrast to optimization-oriented methods, the images condensed by M3D retain more natural features and are more visually recognizable to humans. More visualization results are provided in the Appendix.
Figure 5: Representative samples condensed by M3D on ImageNet. The corresponding labels, from left to right and top to bottom, are bonnet, green snake, langur, doberman, gyromitra, saluki, vacuum, window screen, and cockroach.